OpenStudy (anonymous):

multiply and simplify(-5/12)(1/4)(-3/7)

OpenStudy (anonymous):

\[(\frac{ -5 }{ 12 })(\frac{1}{4})(\frac{-3}{7}) = \frac{ (-5)(1)(-3) }{ (12)(4)(7) }\]

OpenStudy (anonymous):

Just multiply everything on the top, multiply everything on the bottom, and then if reduce the fraction if you can.

OpenStudy (anonymous):

do I multiply them 1 by 1

OpenStudy (anonymous):

Sure, on top, you have (-5) x (1) x (-3) So, (-5) x (1) is just (-5).... then (-5) x (-3) = 15 Can you do the bottom?

OpenStudy (anonymous):

nope im lost #sorry

OpenStudy (anonymous):

is it 15/336

OpenStudy (anonymous):

yes, that's right... but you can reduce it further... divide the top and bottom by 3.

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

15 / 336 = (15/3) / (336/3) divide numerator and denominator by 3 = 5 / 112 It's just like you can reduce 2/8 by realizing it's the same as 1/4... you divide top and bottom by 2 to get 1/4. Here, both numerator and denominator can be divided evenly by 3. 15 5 x 3 5 -- = ------- = -- 336 112 x 3 112

OpenStudy (anonymous):

how did you know it was by 3

OpenStudy (anonymous):

Well, I wanted to reduce it if possible... it looked "too big"... so I knew the factors of 15 were 3 and 5. When I looked at 336, I knew it wasn't divisible by 5, since it ends in 6. I wasn't sure 336 was divisible by 3, but I tried it, and it worked evenly.

OpenStudy (anonymous):

If I had tried 3 and 5 and neither factor would divide evenly into 336, we would have just needed to leave it as "15/336" You can't always reduce, but if you can, you should.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome... hope it helped :)